Luciano wants to have a maximum of $120 cash at the ticket booth when his church carnival opens. He will have $1 bills and $5 bills. If x is the number of $1 bills and y is the number of $5 bills the inequality x+5<=120 models the situation. Graph the inequality
List three solutions to the inequality x+5<=120 where both x and y are integers.
First Solution= (
,
)
Second Solution= (
,
)
Third Solution= (
,
)

Answer by math_tutor2020(3817)


x = number of $1 bills = value of said bills
y = number of $5 bills
5y = value of just the $5 bills only
x+5y = total value
x+5y <= 120 is the inequality to set up. You forgot to put a "y" after the 5.

Let's say x = 5.
Let's determine what y could be based on it.
x+5y <= 120
5+5y <= 120
5y <= 120-5
5y <= 115
y <= 115/5
y <= 23
If you had 5 one dollar bills, then you could have at most 23 $5 bills

So one solution could be (5,20)
Check:
x+5y <= 120
5+5*20 <= 120
5+100 <= 120
105 <= 120
The true statement at the end confirms this solution.

The point (5,20) is in the shaded region of the graph.
The graph consists of a straight solid line through (0,24) and (120,0). Points on the solid boundary are solutions. The shaded region is below this solid boundary line.
I'll let the student set up the graph and determine other solutions.